# Imperceptible Protection against Style Imitation from Diffusion Models

### v2

@@ Introduction
Difficulty-Aware Protection
We observed that the difficulty of protection varies across images (or artworks); some images are protected properly with low protection strength, while others require higher strength to achieve the same protection level. Figure 4 shows the protection efficacy of PhotoGuard (Salman et al. 2023) in 20 artworks. The score is calculated as $(\text{FID}_{\text{protected}}-\text{FID}_{\text{unprotected}})/\text{FID}_{\text{unprotected}}$, with a higher value indicating better protection efficacy. Surprisingly, the protection efficacy varies greatly across artworks. Moreover, some artworks show no difference between unprotected and protected cases. It implies that, for actual users, the average score might be a meaningless metric. So why does the level of protection differ between artworks? This question may tie into the learning dynamics of personalization methods, which are still unclear and require further investigation.
Nevertheless, we understand that the protection performance differs between artworks, so in practice, we can adjust the protection strength accordingly when we know the difficulty. The most straightforward way to predict the difficulty is to perform protection, personalize the protected/unprotected images with DreamBooth (Ruiz et al. 2023), and evaluate protection efficacy as done in Figure 4. While effective, this is extremely time-consuming as it requires two personalization stages (for protected and unprotected images).
Table 1: Quantitative comparison of protection methods w/ and w/o Impasto, both selected for their comparable protection performance. Impasto markedly elevates the protected images’ quality while maintaining protection efficacy.
@@
To predict difficulty more practically, we approximate the above procedure. At the $i$-th perturbation optimization step, we obtain an intermediate protected image as $\mathbf{x}^{(i)}=\mathbf{x}+\delta^{(i)}$. Then, along with the unprotected image $\mathbf{x}$, we create a difficulty map $\mathcal{M}_{D}$ with a difficulty-aware protection function; $\mathcal{M}_{D}=\text{DAP}(\mathbf{x},\,\mathbf{x}^{(i)})$. This function is defined as follows:
$\text{DAP}(\mathbf{x},\,\mathbf{x}^{(i)})=\text{LPIPS}(\text{DP}(\mathbf{x}),\,\text{DP}(\mathbf{x}^{(i)})),$ (8)
where $\text{LPIPS}(\cdot)$ denotes the spatial LPIPS distance (Zhang et al. 2018). In the difficulty prediction module, denoted as $\text{DP}(\cdot)$, an input image $\mathbf{x}$ (or the unprotected image $\mathbf{x}^{(i)}$ but omitting the superscript for clarity) is passed through an encoder $\mathcal{E}$ to obtain the latent code $\mathbf{z}$. A forward diffusion process is then applied for $t$ timestep, followed by a reverse diffusion process using the denoising UNet, $\epsilon_{\theta}$. Finally, the decoder $\mathcal{D}$ projects this back to the image space:
@@
$\tilde{\mathbf{z}}_{0}=\text{Reverse}(\mathbf{z}_{t};\,\epsilon_{\theta}),\,\,\mathbf{z}_{t}=\text{Forward}(\mathcal{E}(\mathbf{x}),\,t),$
where $\text{Forward}(\cdot)$ and $\text{Reverse}(\cdot)$ denote the forward and reverse diffusion processes, respectively. We limit the diffusion process to $t$ timestep to reduce major structural changes.
The difficulty and the perceptual map are used to compute the protection image as: $\hat{\mathbf{x}}=\mathbf{x}+\mathcal{M}\odot\mathcal{M}_{D}\odot\delta$. In this approximated approach, we can replace the difficulty prediction involving full personalization with the diffusion process and LPIPS computation. While full prediction is optimal, we make this approximation to balance practicality. We observe that images with our difficulty-aware protection performed robustly even without the personalization process.
Perceptual Constraint Bank
@@
Masked LPIPS. LPIPS (Zhang et al. 2018) is a widely used constraint. Our approach distinguishes itself by applying a masked LPIPS constraint, modulating the LPIPS influence using a perceptual map $\mathcal{M}$. Let $\phi_{l}$ be a $l$-th layer of the LPIPS network, with a corresponding feature map resolution $d_{l}$, the masked LPIPS is calculated as:
$\mathcal{L_{L}}=\sum_{l}\frac{1}{d_{l}}\sum^{d_{l}}_{i=1}\mathcal{M}_{i}*||w_{l}*(\phi_{l}(\mathbf{x})_{i}-\phi_{l}(\mathbf{x}+\delta)_{i})||^{2}_{2},$ (10)
@@
Masked low-pass. We also apply a pixel-domain constraint that focuses on the low-frequency components, inspired by Luo et al. (2022). The loss function is as:
$\mathcal{L_{LP}}=\frac{1}{d}\sum^{d}_{i=1}\mathcal{M}_{i}*||\text{LP}(\mathbf{x})_{i}-\text{LP}(\mathbf{x}+\delta)_{i}||^{2}_{2},$ (11)
where $\text{LP}(\mathbf{x})$ is the low-frequency component of image $\mathbf{x}$ (please refer to Suppl. for more details). This constraint mimics observing a painting from a distance, where perturbations in smooth regions are perceptible, while detailed textures hide these until we closely inspect the artwork.
Table 2: Component analysis. PAP: perception-aware protection. w/o JND: initialize $\mathbf{M}$ with random masks, not from JNDs. w/o Mask: Constraints without a perceptual map. DAP: difficulty-aware protection. Base*: PhotoGuard with equal protection strength. Base**: PhotoGuard with a similar protection performance (lower protection strength).
@@
CLIP. We leverage CLIP (Radford et al. 2021) which benefits from training on a vast and varied dataset of image-text pairs. With the prompt $C=$ “Noise-free image”, the CLIP constraint aims to maximize the feature distance between the protected image and the descriptive prompt.
where $\text{CLIP}_{I},\text{CLIP}_{T}$ are image and text encoders. The final protection loss, $\mathcal{L}(\mathbf{x}^{(i)},\delta,\mathbf{y},\mathcal{M})$ combines all the losses, weighted by their respective $\lambda$s as:
Algorithm 1 overviews the protection process of Impasto. It is designed to be versatile, allowing the integration of existing protection frameworks. Such adaptability will be discussed in the next section as well as in Suppl.
Experiment
Datasets. We utilize two art domain datasets: painting and cartoon. The painting dataset is curated from WikiArt (Tan et al. 2018) with a selection of 15 artists, 10 works per artist. The cartoon dataset is a collection of 15 cartoons with 10 cartoon face images. Further details are in Suppl.
Table 3: Perception-aware protection. The efficacy of JNDs (LA and CM as they show best results) is presented along with an averaged perceptual map and with IWR. IWR: IWR is conducted before optimization.
@@
(a) Painting dataset
(b) Cartoon dataset
Measure. We use DISTS (Ding et al. 2020), PieAPP (Prashnani et al. 2018) and TOPIQ (Chen et al. 2023) to assess protected image quality. Protection score is measured via NIQE (Mittal, Soundararajan, and Bovik 2012), BRISQUE (Mittal, Moorthy, and Bovik 2011), and FID (Heusel et al. 2017). For protection metrics, worse scores indicate more effective protection.
Baseline. Impasto can be integrated into any existing protection methods. Hence, we adopt Impasto into PhotoGuard (Salman et al. 2023) and AdvDM (Liang et al. 2023). In Suppl, applications to Mist (Liang and Wu 2023) and Anti-DreamBooth (Van Le et al. 2023) are also presented.
Model Comparison
Table 1 presents a comparison of image quality under a comparable protection performance of the models with and without Impasto. Across all the scenarios, Impasto substantially enhances the fidelity of the protected images. Figure 5 also supports the superior efficacy of Impasto; it successfully minimizes artifacts, in contrast to baselines that leave discernible traces on the artwork. User evaluation further confirms the effectiveness of Impasto (Figure 6).
@@
Varying strengths. As demonstrated in Figure 7, we manipulate the protection strengths (budget) to delineate an imperceptibility-protection trade-off curve. On both painting and cartoon datasets, employing Impasto into the baselines considerably improves the trade-off dynamics.
Model Analysis
@@
Ablation study. As shown in Table 2, the perception-aware protection (PAP) markedly improves image fidelity over Photoguard (Base*), albeit with a slight reduction in protection efficacy. However, compared to Photoguard with lower strength (Base**), PAP achieves better preservation of protection performance with enhanced image quality. We observed that when $\mathcal{M}$ is formed from a random mask, not JND, (w/o JND), the protection performance is degraded. Employing a constraint bank significantly improves image quality with comparable protection scores to Photoguard (Base**). The omission of the perceptual map $\mathcal{M}$ in the constraints (w/o Mask) leads to a decline in protection performance, akin to the observations in the PAP case. The difficulty-aware protection (DAP) also enhances both image quality and protection efficacy (FID).
@@
Perceptual map. We analyze the perception-aware protection in Table 3. LA and CM enhance the protected image’s fidelity but at the cost of compromising protection performance. Multiple JNDs through averaging leads to better image quality, but it also reduces the protection performance. We speculate that such a straightforward averaging method may not adequately capture the unique structural elements of the image, resulting in overly smooth perturbations that could weaken the protection performance. On the other hand, IWR enhances all protection scores while preserving satisfactory image quality, as it can adapt to the specific textures and structures of a given artwork. It’s noteworthy that applying IWR prior to perturbation optimization (IWR), where it does not consider the perturbations, slightly diminishes protection performance, accentuating the importance of modeling the interplay between artwork and applied perturbation to finalization mask $\mathcal{M}$.
Difficulty-aware protection. In Figure 8, we demonstrate the change in protection performance with and without difficulty-aware protection (DAP). This method applies stronger intensity in cases of higher protection difficulty and vice versa, leading to a more balanced protection trend.
Countermeasures. To analyze the robustness of Impasto, we conduct countermeasure experiments with JPEG compression ($q=40$), Gaussian blur ($3\times 3$ kernel, $\sigma=0.02$), and Gaussian noise ($\sigma=0.02$). Results indicate a performance degradation of all protection methods when these countermeasures are applied, as they tend to remove the protective perturbations (Figure 9). Nonetheless, Impasto demonstrates comparable robustness against such countermeasures. For the blur effect, there is a slight performance degradation with Impasto, as subtle perturbations are particularly weakened to this; we also observed that baseline with low-budget protection is also vulnerable to blur. However, the blurring effect also renders the artwork less plausible, making it an impractical choice for malicious users.

### v4

@@ III-B Difficulty-Aware Protection
TABLE I: Quantitative comparison of protection methods w/ and w/o Impasto, both selected for their comparable protection performance. Impasto markedly elevates the protected images’ quality while maintaining protection efficacy.
@@
Anti-DB 0.260 (+0.000) 0.154 (+0.000) 0.700(+0.000) 4.437 15.48 160.6
Mist 0.256 (+0.000) 0.160 (+0.000) 0.709(+0.000) 4.597 10.86 158.7
Diff-Protect 0.258 (+0.000) 0.236 (+0.000) 0.758(+0.000) 4.704 13.27 159.2
+ Impasto 0.236 (+0.022) 0.145 (+0.091) 0.838(+0.080) 4.792 12.28 160.1
Figure 6: Qualitative comparison. We visualize the difference map ($\Delta$) between the protected and unprotected images. While maintaining comparable style protection efficacy (artifacts in the generated results), Impasto significantly enhances the quality of the protected images.
Approach. Based on these observations, we propose an approximated approach for difficulty-aware protection (DAP) that avoids the full personalization. In essence, we estimate how easily the introduced protective perturbation can be “undone” by a partial diffusion process [35], and use this estimate as a proxy for the protection difficulty of the image.
Specifically, given the original image $\mathbf{x}$ and the $i$-th protected image $\mathbf{x}^{(i)}=\mathbf{x}+\delta^{(i)}$, we first extract latent codes $\mathbf{z}$ and $\mathbf{z}^{(i)}$ using an encoder $\mathcal{E}$. We then apply $t$ steps of forward diffusion to each latent and perform $t$ steps of reverse denoising using a diffusion UNet $\epsilon_{\theta}$ [35]. Finally, the denoised latents are projected back to the image space using a decoder $\mathcal{D}$, yielding reconstructed images $\text{DP}(\mathbf{x})$ and $\text{DP}(\mathbf{x}^{(i)})$. This process is illustrated in Eq. 8.
@@
$\tilde{\mathbf{z}}_{0}=\text{Reverse}(\mathbf{z}_{t};\,\epsilon_{\theta}),\,\,\mathbf{z}_{t}=\text{Forward}(\mathcal{E}(\mathbf{x}),\,t).$
Here, $\text{Forward}(\cdot)$ and $\text{Reverse}(\cdot)$ denote the forward and reverse diffusion processes, respectively.
Intuitively, if the reconstructed image $\text{DP}(\mathbf{x}^{(i)})$ retains a large amount of visible perturbation, it suggests that the original image $\mathbf{x}$ is relatively easy to protect. In contrast, if $\text{DP}(\mathbf{x}^{(i)})$ closely resembles $\text{DP}(\mathbf{x})$, it indicates that the protection is more difficult, as the noise was effectively reversed. Note that we set $t=5$ out of $T=25$ (total denoising steps), to ensure that the diffusion process captures local perturbations without inducing major scene changes. This partial diffusion is computationally efficient and sufficiently approximates how DreamBooth may leverage or undo the added noise.
To quantify the difficulty, we compute the spatial LPIPS distance [70] between the two outputs. LPIPS is chosen due to its strong correlation with perceptual differences in images.
$\text{DAP}(\mathbf{x},\mathbf{x}^{(i)})=\text{LPIPS}\bigl(\text{DP}(\mathbf{x}),\,\text{DP}(\mathbf{x}^{(i)})\bigr).$ (9)
Consequently, $\text{DAP}(\mathbf{x},\mathbf{x}^{(i)})$ acts as a spatial difficulty map, which is multiplied element-wise with the perceptual map $\mathcal{M}$ to selectively reinforce protection in more vulnerable regions:
$\hat{\mathbf{x}}=\mathbf{x}+\delta\,\odot\,\mathcal{M}\,\odot\,\underbrace{\text{DAP}(\mathbf{x},\mathbf{x}^{(i)})}_{\mathcal{M}_{D}}.$ (10)
In our DAP, when $DP(\cdot)$ significantly alters the protected image compared to the original (i.e. yields a high LPIPS distance), it implies that the protection is easily neutralized by the diffusion models, thus requires stronger protection. Conversely, a low LPIPS distance suggests that the protection is resilient thus the region requires less protection strength, as personalization is less likely to succeed. While our process does not fully replicate DreamBooth’s behavior, our empirical results show a strong correlation between the proposed surrogate measure and actual personalization performance. This enables effective, difficulty-aware protection without the computational overhead of full DreamBooth tuning.
III-C Perceptual Constraint Bank
@@
Masked LPIPS. LPIPS [70] is a widely used constraint. Our approach distinguishes itself by applying a masked LPIPS constraint, modulating the LPIPS influence using a perceptual map. Let $\phi_{l}$ be a $l$-th layer of the LPIPS network, with a corresponding feature map resolution $d_{l}$, the masked LPIPS is calculated as in below equation.
[…]
@@
Masked low-pass. We also apply a pixel-domain constraint that focuses on the low-frequency components, inspired by Luo et al. [32]. The loss function is as:
[…]
where $\text{LP}(\mathbf{x})$ is the low-frequency component of image $\mathbf{x}$ (please refer to Supplemental Materials for more details). This constraint mimics observing a painting from a distance, where perturbations in smooth regions are perceptible, while detailed textures hide these until we closely inspect the artwork.
CLIP. We leverage CLIP [46] which benefits from training on a vast and varied dataset of image-text pairs. With the prompt $C=$ “Noise-free image”, the CLIP constraint aims to maximize the feature distance between the protected image and the descriptive prompt.
[…]
The perceptual constraint bank works as an ensemble of multi-feature constraints. While prior studies in adversarial attacks or image protection use constraints, they typically rely on only one or two types. Interestingly, we argue that applying constraints across multiple feature spaces can significantly enhance performance, as each feature space captures distinct aspects of the image. Despite its potential, such an approach has not been explored in previously. Algorithm 1 overviews the protection process of Impasto. It is designed to be versatile, allowing the integration of existing protection frameworks.
III-D Implementation Details
When we implement Impasto into the existing protection methods, we adhere to their respective settings. Regarding the Impasto’s own components, we use following hyperparameters depicted in Eq. 14 as: $\lambda_{L}=5.0,\lambda_{LP}=10.0,\lambda_{C}=0.1$. For our PAP (Eq. 6), we utilize the $L_{\infty}$ norm. On the other hand, for the IWR loss (Eq. 7), we employ the $L_{2}$ norm. When we calculate IWR loss, since the distance between $\mathcal{L_{SP}}$ respect to $\mathcal{M}^{\prime}$ and $\mathcal{M}(\omega)$ (first term) is significantly smaller than the $L_{2}$ constraint (second term), we amplify the former term by a factor of $5\times 10^{7}$. For difficulty prediction in DAP, we set $t=5$ with $T=25$. We execute IWR and DAP every $P=4$ steps. On an A100 GPU, the overall perturbation optimization takes approximately 30 seconds per $\times512$ image when Impasto is integrated with PhotoGuard. In comparison, the original PhotoGuard requires around 10 seconds under the same environment. Although our method is relatively slower, we prioritize imperceptibility; a practically critical aspect for real-world usage.
In addition, during our experiments, we observed that the magnitudes of $\mathcal{L_{E}}$ and $\mathcal{L_{SD}}$ differ considerably, with the SD loss exhibiting much smaller values. Hence, when integrating Impasto into UNet-based protection methods (e.g. AdvDM [29], Anti-DreamBooth [59], Mist [30]), we scale down $\lambda_{L},\lambda_{LP},\lambda_{C}$ by a factor of 0.05. This adjustment is made to balance the influence of our proposed components, ensuring that the impact of Impasto is consistent and effective in enhancing image protection. Last, the experiments in this paper were conducted based on SD v1.5 [47], utilizing DreamBooth [48] as the method of personalization. The generalization performance of the diffusion model and personalization is provided separately in the experimental section.
IV Experiment
Datasets. We utilize two art domain datasets: painting and cartoon. The painting dataset is curated from WikiArt [58] with a selection of 15 artists, 10 works per artist. The cartoon dataset is a collection of 15 cartoons with 10 cartoon face images. Further details are in Supplemental Material.
Measure. We use DISTS [10], PieAPP [45] and TOPIQ [4] to assess protected image quality. Protection score is measured via NIQE [38], BRISQUE [37], and FID [16]. For protection metrics, higher scores indicate stronger protection (e.g. a higher FID implies better protection performance).
TABLE II: Component analysis. PAP: perception-aware protection. w/o JND: initialize $\mathbf{M}$ with random masks, not from JNDs. w/o Mask: Constraints without a perceptual map. DAP: difficulty-aware protection. Base*: PhotoGuard with equal protection strength. Base**: PhotoGuard with a similar protection performance (lower protection strength).
@@ IV Experiment
TABLE III: Component analysis. Within Impasto, we conduct an analysis by excluding the JND-based perceptual map generation, resulting in the creation of the perceptual map from random masks (denoted as w/o JND). Additionally, we evaluate the impact of omitting the perception-aware protection (denoted as w/o PAP).
Method Protected Image Quality Protection Performance
Impasto 0.159 0.912 20.74 279.3
w/o JND 0.158 0.912 19.30 268.1
w/o PAP 0.199 0.890 17.57 281.1
User study. We conducted a user study in the form of an A/B test with a reference image as a benchmark. A total of 60 participants were involved, tasked with determining the better-quality protected image (given the reference original artwork) and identifying the lower-quality generated image (given a reference DreamBooth [48] generated image). Each participant was asked to vote on 8 questions for PhotoGuard [50] and another 9 for AdvDM [29].
Baseline. Impasto can be incorporated into any existing protection method. Therefore, we apply Impasto to both encoder and diffusion-based protection frameworks. In our model comparison, we evaluate the performance of methods with and without Impasto, using PhotoGuard [50], AdvDM [29], Mist [30], Diff-Protect [67], and Anti-DreamBooth [59] as baselines.
IV-A Model Comparison
Prior work compares protection methods under a fixed perturbation strength. However, we observe that both protection efficacy and imperceptibility can vary considerably depending on the specific loss function employed. For example, PhotoGuard [50] and AdvDM [29] follow distinct optimization objectives, leading to different trade-off patterns between visibility and robustness. Furthermore, focusing solely on protection performance without considering visual imperceptibility provides an incomplete picture that may not align with real-world deployment needs. To account for this, we adaptively adjust the perturbation strength for each method to achieve comparable protection levels (similar protection efficacy).
Table I presents a comparison of image quality under a comparable protection performance of the models with and without Impasto. Notably, Impasto achieves comparable or even superior protection strength while significantly improving visual quality, thereby offering a more favorable trade-off between robustness and invisibility. Across all the scenarios, Impasto substantially enhances the fidelity of the protected images. Fig. 15 also supports the superior efficacy of Impasto; it successfully minimizes artifacts, in contrast to baselines that leave discernible traces on the artwork. It is particularly pronounced in the cartoon dataset (bottom), where both vanilla methods introduce noticeable artifacts in facial areas, potentially disrupting user immersion when reading cartoons. In contrast, Impasto reduces artifacts significantly, rendering them nearly invisible unless examined closely and meticulously. This suggests that our method ensures stronger protection with minimal perceptual degradation, which is especially beneficial for sensitive visual domains such as character-centric content. User evaluation further confirms the effectiveness of Impasto (Fig. 7).
TABLE IV: Perception-aware protection. The efficacy of JNDs (LA and CM as they show best results) is presented along with an averaged perceptual map and with IWR. IWR: IWR is conducted before optimization.
@@ IV-A Model Comparison
Varying strengths. As demonstrated in Fig. 8, we manipulate the protection strengths (budget) to delineate an imperceptibility-protection trade-off curve. On both painting and cartoon datasets, employing Impasto into the baselines considerably improves the trade-off dynamics. Specifically, Impasto consistently yields higher image quality at the same protection level, or stronger protection at the same perceptual budget, highlighting a more favorable imperceptibility-robustness balance compared to existing methods.
IV-B Model Analysis
@@
Ablation study. As shown in Table II, the perception-aware protection (PAP) markedly improves image fidelity over Photoguard (Base*), albeit with a slight reduction in protection efficacy. However, compared to Photoguard with lower strength (Base**), PAP achieves better preservation of protection performance with enhanced image quality. We observed that when $\mathcal{M}$ is formed from a random mask, not JND, (w/o JND), the protection performance is degraded. Employing a constraint bank that combines LPIPS, low-pass, and CLIP constraints significantly improves image quality while maintaining protection performance comparable to Photoguard (Base**). LPIPS emphasizes perceptual similarity, the low-pass filter suppresses high-frequency noise, and CLIP ensures semantic consistency. These components work synergistically to balance visual refinement and robust protection, validating the method’s effectiveness. The omission of the perceptual map $\mathcal{M}$ in the constraints (w/o Mask) leads to a decline in protection performance, akin to the observations in the PAP case. The difficulty-aware protection (DAP) also enhances both image quality and protection efficacy (FID).
One might expect that PAP with JND maintains protection scores while improving image quality. Indeed, the random mask (PAP without JND) constrains the protection intensity in a similar level to the JND-based map. This is because both maps are normalized between 0 and 1. As a result, they yield similar perturbation magnitudes, leading to a comparable image quality. Nonetheless, the JND-based perceptual map prioritizes less sensitive regions for perturbation while reducing it in highly sensitive areas. Despite the similar protection magnitudes, the JND-based PAP achieves more effective protection. This hints at the possibility that the JND maps are helpful in finding the areas that are important for style protection. For instance, in areas with complex textures, the PAP applies more perturbations than its non-JND counterpart, yet these perturbations remain imperceptible to humans, thereby ensuring perceptually acceptable image quality. Moreover, the interaction between the JND-based perceptual map and PAP drives a synergistic effect, where the former guides the spatial distribution of perturbations and the latter enforces robust protection, collectively enhancing the overall protection efficacy.
In Table III, we show additional analysis by omitting certain components from the full Impasto framework. Specifically, we assess the impact of 1) removing the JND-based perception map, which results in initializing $\mathbf{M}$ with random masks (w/o JND), and 2) excluding the proposed perception-aware protection (PAP), thereby relying solely on the perceptual constraint bank (w/o PAP). For the w/o JND case, in contrast to Table II, we start with a full-component model and remove only the JND part. Results indicate that discarding the JNDs leads to a decreased protection performance with a similar image quality, corroborating the findings in Table II. When PAP is not applied, there is a slight increase in protection performance (FID: 279.3 $\rightarrow$ 281.1) with significant compromise on image quality. These observations highlight that the combined interaction of PAP and the JND-based perceptual map is instrumental in achieving an imperceptible style protection framework by efficiently balancing high protection performance with superior image quality.
(a) PhotoGuard + Impasto
(b) AdvDM + Impasto
@@
JNDs. Here, we further discuss the protection performance of JND-based perception map depicted in Fig. 3 (right). When compared to the baseline, perceptual map improves image quality (DISTS) across the board, but it also leads to a compromise in protection performance (FID). Among the JNDs, LA demonstrates the best protection performance. We conjecture that in many artworks, the majority of areas fall high or low-luminance, thereby maintaining perturbations strength high across extensive regions. However, as shown in upper image, even simple textures (e.g. sky) can have strong perturbations, placing LA at the lower image quality. On the other hand, CSF, Std, and Entropy generally apply high perturbations only to specific areas, such as edges, resulting in most regions being not fully protected and consequently, causing a huge degradation in protection performance. CM’s protection intensity is also determined by spatial changes but this covers more detailed local regions (see fields in upper image) and also being based on contrast, leading to both high quality and effective protection.
Perceptual map. We analyze the perception-aware protection in Table IV. LA and CM enhance the protected image’s fidelity but at the cost of compromising protection performance. Multiple JNDs through averaging leads to better image quality, but it also reduces the protection performance. We speculate that such a straightforward averaging method may not adequately capture the unique structural elements of the image, resulting in overly smooth perturbations that could weaken the protection performance. On the other hand, IWR enhances all protection scores while preserving satisfactory image quality, as it can adapt to the specific textures and structures of a given artwork. It’s noteworthy that applying IWR prior to perturbation optimization (IWR), where it does not consider the perturbations, slightly diminishes protection performance, accentuating the importance of modeling the interplay between artwork and applied perturbation to finalization mask $\mathcal{M}$.